OpenStudy (anonymous):

Integrate (1+x^2)/(1+x^4)dx

OpenStudy (anonymous):

\[\large \int\limits\frac{1+x^{2}}{1+x^{4}} dx\]

OpenStudy (lgbasallote):

ugh why couldnt it be 1 - x^4 :p lol

OpenStudy (anonymous):

If it was it would have defeated the purpose of this integral. (:

OpenStudy (lgbasallote):

exactly why im complaining :p

OpenStudy (anonymous):

No need to complain. Just think of an intuitive method. :)

OpenStudy (blockcolder):

Completing the square of the denominator then a substitution is what I have in mind, but I'm not completely sure.

OpenStudy (anonymous):

Not that easy.

OpenStudy (anonymous):

There are two methods.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integrate+%5Cfrac%7B1%2Bx%5E%7B2%7D%7D%7B1%2Bx%5E%7B4%7D%7D+dx (look at "Show steps") Well, this looks... difficult. I wonder if there's an easier way to do this, than splitting it into two integrals.

OpenStudy (lgbasallote):

okay...hmm \[u = \frac{1}{1+x^4} \rightarrow du = \frac{4x^3}{(1+x^4)^2}\] \[dv = 1+ x^2 \rightarrow v = \tan^{-!} x\] LOL really isn't easy :P

OpenStudy (blockcolder):

Separation, perhaps? \[\frac{1}{1+x^4}+\frac{x^2}{1+x^4}\] Might this work?

OpenStudy (anonymous):

Not really. Wolfram Alpha method is tedious.

OpenStudy (anonymous):

You can try if it works; but I dont think so.

OpenStudy (lgbasallote):

the first term would be integrable i think

OpenStudy (anonymous):

\[\large\frac{1+\frac1{x^2}}{x^2 + \frac{1}{x^2}} = \frac{1 + \frac1{x^2}}{\left(x -\frac{1}x\right)+2} \]This should do it.

OpenStudy (blockcolder):

Wow. O_O

OpenStudy (anonymous):

\[\large\frac{1+\frac1{x^2}}{x^2 + \frac{1}{x^2}} = \frac{1 + \frac1{x^2}}{\left(x -\frac{1}x\right)^2+2}\]

OpenStudy (anonymous):

Very nice Ishaan94; trig sub would work as well (: But that is easier.